During an exam, a vet found that only 3/5 of an animals teeth had emerged. the assistant counted 15 teeth how many will the animal eventually have?
Answer by rfer(16322)

15/.6=25
or
(15/1)/(3/5)
(15/1)(5/3)=75/3=25